# Quantitative convergence of trained single layer neural networks to Gaussian processes

### v2

@@ 3 Assumptions and main result
Let $f^{\textup{lin}}$ be the linearization of $f$, and let $x\in\mathbb{R}^{n_{0}}$ be a test point. Then, under assumptions 1 and 3, there exist positive constants $C,\overline{C},\overline{D}$ not depending on $n_{1}$ nor $t$ such that:
$\mathcal{W}_{2}^{2}(f^{\textup{lin}}(\mathcal{X};\overline{\theta}_{t}),G_{t}(\mathcal{X}))$ $\leq\frac{1}{n_{1}}C(t+1)e^{-\lambda_{\textup{min}}^{\infty}t},$
$\mathcal{W}_{2}^{2}(f^{\textup{lin}}(x;\overline{\theta}_{t}),G_{t}(x))$ $\leq\frac{1}{n_{1}}(\overline{C}+\overline{D}(t+1)e^{-\lambda_{\textup{min}}^{\infty}t}).$
@@ B.1 Results at initialization
$\mathbb{E}[|\tilde{k}_{11}-\mathcal{K}|^{p}]$ $=0,$ (B.7)
@@
Note that since $k_{\infty}$ is deterministic, we have $\mathcal{W}_{p}^{p}(k,k_{\infty})=\mathbb{E}[\|k-k_{\infty}\|^{p}]$. The constant $C$ in B.4 depends on the constants produced by applications of Theorem B.3.
@@ B.2 Approximation by linearization at training time $t>0$
Fix $x\in\mathbb{R}^{n_{0}}$ a test point. Let $G_{t}^{x}=G_{t}(x)$ and $G_{t}^{\mathcal{X}}=G_{t}(\mathcal{X})$.
First we show the result for the training set. With the aid of Equation (2.2), we derive a closed ODE for $|f_{t}-G_{t}|^{2}$. By Cauchy-Schwarz’s inquality and Young’s inequality
@@
Note that, by gradient flow equation we have $\|f_{t}-y\|\leq e^{-\frac{\lambda_{\textup{min}}^{\infty}}{2}t}\|f_{0}-y\|$ for each $t\geq 0$. Choosing ${\varepsilon}=\lambda_{\textup{min}}^{\infty}$ and putting $b_{t}=\frac{e^{-\lambda_{\textup{min}}^{\infty}t}}{\lambda_{\textup{min}}^{\infty}}\|k_{\mathcal{X}\mathcal{X}}-k_{\infty}\|^{2}\|y-f_{0}\|^{2}$ yields:
$\frac{\partial}{\partial t}\|f_{t}-G_{t}^{\mathcal{X}}\|^{2}\leq b_{t}-\lambda_{\textup{min}}^{\infty}\|f_{t}-G_{t}^{\mathcal{X}}\|^{2}.$ (D.5)
@@
$\|f_{t}-G_{t}\|^{2}$ $\leq e^{-\lambda_{\textup{min}}^{\infty}t}\left(\|f_{0}-G_{0}\|^{2}+\int_{0}^{t}e^{\lambda_{\textup{min}}^{\infty}s}b_{s}ds\right)$ (D.6)
$=e^{-\lambda_{\textup{min}}^{\infty}t}\left(\frac{\|k_{\mathcal{X}\mathcal{X}}-k_{\infty}\|^{2}\|f_{0}-y\|^{2}t}{\lambda_{\textup{min}}^{\infty}}+\|f_{0}-G_{0}\|^{2}\right).$ (D.7)
Recall the definition of $2$-Wasserstein distance. Taking the expected value of the previous equation and taking the infimum on all the couplings between $f_{t}$ and $G_{t}$ we can bound by Hölder’s inequality:
@@
$\leq e^{-\lambda_{\textup{min}}^{\infty}t}\left(\frac{t}{\lambda_{\textup{min}}^{\infty}}\mathbb{E}[\|k_{\mathcal{X}\mathcal{X}}-k_{\infty}\|^{2}\|f_{0}-y\|^{2}]+\mathbb{E}[\|f_{0}-G_{0}\|^{2}]\right)$ (D.9)
$\leq e^{-\lambda_{\textup{min}}^{\infty}t}\left(\frac{t}{\lambda_{\textup{min}}^{\infty}}\mathbb{E}[\|k_{\mathcal{X}\mathcal{X}}-k_{\infty}\|^{4}]^{\frac{1}{2}}\mathbb{E}[\|f_{0}-y\|^{4}]^{\frac{1}{2}}+\mathbb{E}[\|f_{0}-G_{0}\|^{2}]\right).$ (D.10)
Now we can take the infimum on the couplings between $f_{0}$ and $G_{0}$ to apply Theorem B.3, Proposition B.4 and Lemma B.8 to estimate the right-hand side of (D.10). There are positive constants $c_{1}$ and $c_{2}$ not depending on $n_{1}$ such that:
$\mathcal{W}^{2}_{2}(f_{t},G_{t})$ $\leq e^{-\lambda_{\textup{min}}^{\infty}t}\left(\frac{c_{1}t}{\lambda_{\textup{min}}^{\infty}n_{1}}\sqrt{32n^{2}\|\Phi\|_{\infty}^{4}+8\|y\|^{4}}+\frac{c_{2}}{n_{1}}\right)$ (D.11)
$\leq\frac{Ce^{-\lambda_{\textup{min}}^{\infty}t}}{n_{1}}(t+1),$ (D.12)
with $C=\max\{\frac{2c_{1}}{\lambda_{\textup{min}}^{\infty}}\sqrt{8n^{2}\|\Phi\|_{\infty}^{4}+2\|y\|^{4}},c_{2}\}$.
@@
By the formula for the derivative of the product and Cauchy-Schwarz’s inequality, the preceding equation implies:
@@
Put $\overline{\lambda}=\min_{1\leq i\leq n}k_{\infty}(x,x_{i})$. The last summand can be further estimated with the first result in this theorem. There exists a positive constants $C$ not depending on $n_{1}$ such that:
$\frac{\partial}{\partial t}(y_{t}-G_{t}^{x})$ $\leq\|k_{x\mathcal{X}}-k_{\infty}^{x}\|\|y-f_{t}\|+\frac{\overline{\lambda}Ce^{-\frac{\lambda_{\textup{min}}^{\infty}}{2}t}}{\sqrt{n_{1}}}\sqrt{t+1}.$ (D.14)
Again, we use $\|f_{t}-y\|\leq e^{-\frac{\lambda_{\textup{min}}^{\infty}}{2}t}\|f_{0}-y\|$. Integrating we obtain:
$(y_{t}-G_{t}^{x})$ $\leq(y_{0}-G_{0}^{x})+\|k_{x\mathcal{X}}-k_{\infty}^{x}\|\|y-f_{0}\|(1-e^{-\lambda_{\textup{min}}^{\infty}t})+\frac{\overline{\lambda}C}{\sqrt{n_{1}}}\int_{0}^{t}\sqrt{s+1}e^{-\frac{\lambda_{\textup{min}}^{\infty}}{2}s}ds$ (D.15)
$\leq(y_{0}-G_{0}^{x})+\|k_{x\mathcal{X}}-k_{\infty}^{x}\|\|y-f_{0}\|(1-e^{-\lambda_{\textup{min}}^{\infty}t})+\frac{\overline{\lambda}CD}{\sqrt{n_{1}}}(2-\sqrt{t+1}e^{-\frac{\lambda_{\textup{min}}^{\infty}t}{2}}),$ (D.16)
for a positive constant $D$, which explicit computation we now show separately. First we compute an antiderivative of $\sqrt{t+1}e^{-\frac{\lambda_{\textup{min}}^{\infty}}{2}t}$:
$\int\sqrt{s+1}$ $e^{-\frac{\lambda_{\textup{min}}^{\infty}}{2}s}ds$ (D.17)
$=2e^{\frac{\lambda_{\textup{min}}^{\infty}}{2}}\int u^{2}e^{-\frac{\lambda_{\textup{min}}^{\infty}}{2}u^{2}}du$ (D.18)
$=2e^{\frac{\lambda_{\textup{min}}^{\infty}}{2}}\left(-\frac{ue^{-\frac{\lambda_{\textup{min}}^{\infty}}{2}u^{2}}}{\lambda_{\textup{min}}^{\infty}}+\int\frac{ue^{-\frac{\lambda_{\textup{min}}^{\infty}}{2}u^{2}}}{\lambda_{\textup{min}}^{\infty}}du\right)+C$ (D.19)
$=2e^{\frac{\lambda_{\textup{min}}^{\infty}}{2}}\left(-\frac{ue^{-\frac{\lambda_{\textup{min}}^{\infty}}{2}u^{2}}}{\lambda_{\textup{min}}^{\infty}}+\frac{\sqrt{\pi}}{\sqrt{2(\lambda_{\textup{min}}^{\infty})^{3}}}\int\frac{2e^{-v^{2}}}{\sqrt{\pi}}dv\right)+C$ (D.20)
$=2e^{\frac{\lambda_{\textup{min}}^{\infty}}{2}}\left(-\frac{ue^{-\frac{\lambda_{\textup{min}}^{\infty}}{2}u^{2}}}{\lambda_{\textup{min}}^{\infty}}+\frac{\sqrt{\pi}\operatorname{erf}{v}}{\sqrt{2(\lambda_{\textup{min}}^{\infty})^{3}}}\right)+C$ (D.21)
$=2e^{\frac{\lambda_{\textup{min}}^{\infty}}{2}}\left(-\frac{\sqrt{s+1}e^{-\frac{\lambda_{\textup{min}}^{\infty}}{2}(s+1)}}{\lambda_{\textup{min}}^{\infty}}+\frac{\sqrt{\pi}\operatorname{erf}{\left(\frac{\sqrt{\lambda_{\textup{min}}^{\infty}(s+1)}}{\sqrt{2}}\right)}}{\sqrt{2(\lambda_{\textup{min}}^{\infty})^{3}}}\right)+C$ (D.22)
$=-\frac{2\sqrt{s+1}e^{-\frac{\lambda_{\textup{min}}^{\infty}s}{2}}}{\lambda_{\textup{min}}^{\infty}}+\frac{\sqrt{2\pi}e^{\frac{\lambda_{\textup{min}}^{\infty}}{2}}\operatorname{erf}{\left(\frac{\sqrt{\lambda_{\textup{min}}^{\infty}(s+1)}}{\sqrt{2}}\right)}}{(\lambda_{\textup{min}}^{\infty})^{\frac{3}{2}}}+C.$ (D.23)
where in the first step we substituted $u=\sqrt{s+1}$, in the third step we substituted $v=\sqrt{\frac{\lambda_{\textup{min}}^{\infty}}{2}}u$ and in the fourth step we used the definition of Gauss error function $\operatorname{erf}(x)$; and $C$ denotes the integration constant. Therefore,
$\int_{0}^{t}\sqrt{s+1}e^{-\frac{\lambda_{\textup{min}}^{\infty}}{2}s}ds$ $=\frac{2-2\sqrt{t+1}e^{-\frac{\lambda_{\textup{min}}^{\infty}t}{2}}}{\lambda_{\textup{min}}^{\infty}}$ (D.24)
$\quad+\frac{\sqrt{2\pi}e^{\frac{\lambda_{\textup{min}}^{\infty}}{2}}\left(\operatorname{erf}{\left(\frac{\sqrt{\lambda_{\textup{min}}^{\infty}(t+1)}}{\sqrt{2}}\right)}-\operatorname{erf}{\left(\frac{\sqrt{\lambda_{\textup{min}}^{\infty}}}{\sqrt{2}}\right)}\right)}{(\lambda_{\textup{min}}^{\infty})^{\frac{3}{2}}}.$ (D.25)
Since $\operatorname{erf}(t)\in]-1,1[$, we can bound:
$\int_{0}^{t}\sqrt{s+1}e^{-\frac{\lambda_{\textup{min}}^{\infty}}{2}s}ds$ $\leq\frac{2}{\lambda_{\textup{min}}^{\infty}}\left(1-\sqrt{t+1}e^{-\frac{\lambda_{\textup{min}}^{\infty}t}{2}}+\sqrt{\frac{2\pi}{\lambda_{\textup{min}}^{\infty}}}e^{-\frac{\lambda_{\textup{min}}^{\infty}}{2}}\right)$ (D.26)
$\leq D(2-\sqrt{t+1}e^{-\frac{\lambda_{\textup{min}}^{\infty}t}{2}}),$ (D.27)
with $D=\frac{2}{\lambda_{\textup{min}}^{\infty}}\max\{1,\sqrt{\frac{2\pi}{\lambda_{\textup{min}}^{\infty}}}e^{-\frac{\lambda_{\textup{min}}^{\infty}}{2}}\}$.
Turning back to (D.16), we can finish by applying the elementary inequality $(a+b+c)^{2}\leq 3a^{2}+3b^{2}+3c^{2}$ for $a,b,c\geq 1$ and Hölder’s inequality. After that, Theorem B.3, Proposition B.4 and Lemma B.8 can be applied in the same fashion as in the proof of the training case, yielding positive constants $d_{1},d_{2}$ and $d_{3}$ such that:
@@
$\leq 3\mathbb{E}[|y_{0}-G_{0}^{x}|^{2}]+3\mathbb{E}[\|k_{x\mathcal{X}}-k_{\infty}^{x}\|^{4}]^{\frac{1}{2}}\mathbb{E}[\|y-f_{0}\|^{4}]^{\frac{1}{2}}(1-e^{-\lambda_{\textup{min}}^{\infty}t})^{2}$ (D.29)
$\quad+\frac{3(\overline{\lambda}CD)^{2}}{n_{1}}(2-\sqrt{t+1}e^{-\frac{\lambda_{\textup{min}}^{\infty}t}{2}})^{2}$ (D.30)
$\leq\frac{3d_{1}}{n_{1}}+\frac{6d_{2}}{n_{1}}\sqrt{8n^{2}\|\Phi\|_{\infty}^{4}+2\|y\|^{4}}(1+e^{-2\lambda_{\textup{min}}^{\infty}t})$ (D.31)
$\quad+\frac{3(\overline{\lambda}CD)^{2}}{n_{1}}(4+(t+1)e^{-\lambda_{\textup{min}}^{\infty}t}).$ (D.32)
By putting $\overline{C}=\max\{3d_{1},12(\overline{\lambda}CD)^{2}\}$ and $\overline{D}=\max\{6d_{2}\sqrt{8n^{2}\|\Phi\|_{\infty}^{4}+2\|y\|^{4}},3(\overline{\lambda}CD)^{2}\}$ we obtain the thesis.
Note that $C,\overline{C}$ and $\overline{D}$ do not depend neither on $n_{1}$ nor $t$. ∎

### v3

@@ 3 Assumptions and main result
Let $f^{\textup{lin}}$ be the linearization of $f$, and let $x\in\mathbb{R}^{n_{0}}$ be a test point. Then, under assumptions 1 and 3, there exist positive constants $C$ and $C^{\prime}$ not depending on $n_{1}$ nor $t$ such that:
$\mathcal{W}_{2}^{2}(f^{\textup{lin}}(\mathcal{X};\overline{\theta}_{t}),G_{t}(\mathcal{X}))$ $\leq\frac{C}{n_{1}},$
$\mathcal{W}_{2}^{2}(f^{\textup{lin}}(x;\overline{\theta}_{t}),G_{t}(x))$ $\leq\frac{C^{\prime}}{n_{1}}(1+t^{2}).$
@@ 6 Acknowledgements
The authors thank Lucia Celli and Luis Maia for pointing out an issue in the proof of Proposition 3.7.
@@ B.1 Results at initialization
$\tilde{k}_{11}-\tilde{\mathcal{K}}$ $=0,$ (B.7)
@@
Note that since $k_{\infty}$ is deterministic, we have $\mathcal{W}_{p}^{p}(k,k_{\infty})=\mathbb{E}[\|k-k_{\infty}\|^{p}]$. The constant $C$ in Proposition B.4 depends on the constants produced by applications of Theorem B.3.
@@ B.2 Approximation by linearization at training time $t>0$
Fix $x\in\mathbb{R}^{n_{0}}$ a test point. During the proof we will use the notations $f_{t}=f_{t}^{\textup{lin}}(\mathcal{X})$, $y_{t}=f_{t}^{\textup{lin}}(x)$, $G_{t}^{x}=G_{t}(x)$ and $G_{t}^{\mathcal{X}}=G_{t}(\mathcal{X})$. Also, we will denote by $\lambda_{\textup{min}}$ and $\lambda_{\textup{min}}^{\infty}$ the minimum eigenvalues of $k_{\mathcal{X}\mathcal{X}}$ and $k_{\infty}$, respectively. First we show the result for the training set. With the aid of Equation (2.2), we derive a closed ODE for $|f_{t}-G_{t}|^{2}$. By Cauchy-Schwarz’s and Young’s inequalities:
@@
Note that, by gradient flow equation and since $k_{\mathcal{X}\mathcal{X}}$ is semipositive definite, the following rough bound holds for each $t\geq 0$:
$\|f_{t}-y\|\leq e^{-\lambda_{\textup{min}}t}\|f_{0}-y\|\leq\|f_{0}-y\|$
Choosing ${\varepsilon}=\lambda_{\textup{min}}^{\infty}$ and putting $b=\frac{1}{\lambda_{\textup{min}}^{\infty}}\|k_{\mathcal{X}\mathcal{X}}-k_{\infty}\|^{2}\|y-f_{0}\|^{2}$ together yield:
$\frac{\partial}{\partial t}\|f_{t}-G_{t}^{\mathcal{X}}\|^{2}\leq b-\lambda_{\textup{min}}^{\infty}\|f_{t}-G_{t}^{\mathcal{X}}\|^{2}.$ (D.5)
@@
$\|f_{t}-G_{t}\|^{2}$ $\leq e^{-\lambda_{\textup{min}}^{\infty}t}\left(\|f_{0}-G_{0}\|^{2}+b\int_{0}^{t}e^{\lambda_{\textup{min}}^{\infty}s}ds\right).$ (D.6)
Recall the definition of $2$-Wasserstein distance. By taking the expected value of the previous equation and the infimum on all the couplings between $f_{t}$ and $G_{t}$ we can bound by Hölder’s inequality after exchanging the integrals with Fubini-Tonelli theorem:
@@
$\leq e^{-\lambda_{\textup{min}}^{\infty}t}\left(\frac{e^{\lambda_{\textup{min}}^{\infty}t}-1}{{\lambda_{\textup{min}}^{\infty}}^{2}}\mathbb{E}[\|k_{\mathcal{X}\mathcal{X}}-k_{\infty}\|^{2}\|f_{0}-y\|^{2}]+\mathbb{E}[\|f_{0}-G_{0}\|^{2}]\right)$ (D.8)
$\leq\frac{1-e^{-\lambda_{\textup{min}}^{\infty}t}}{{\lambda_{\textup{min}}^{\infty}}^{2}}\mathbb{E}[\|k_{\mathcal{X}\mathcal{X}}-k_{\infty}\|^{4}]^{\frac{1}{2}}\mathbb{E}[\|f_{0}-y\|^{4}]^{\frac{1}{2}}+e^{-\lambda_{\textup{min}}^{\infty}t}\mathbb{E}[\|f_{0}-G_{0}\|^{2}].$ (D.9)
Now we can take the infimum on the couplings between $f_{0}$ and $G_{0}$ on both sides of the inequality to apply Theorem B.3, Proposition B.4 and Lemma B.8 to estimate the right-hand side of (D.9). There are positive constants $c_{1}$ and $c_{2}$ not depending on $n_{1}$ nor $t$ and such that:
$\mathcal{W}^{2}_{2}(f_{t},G_{t})$ $\leq\left(c_{1}\frac{1-e^{-\lambda_{\textup{min}}^{\infty}t}}{{\lambda_{\textup{min}}^{\infty}}^{2}n_{1}}\sqrt{32n^{2}\|\Phi\|_{\infty}^{4}+8\|y\|^{4}}+\frac{c_{2}}{n_{1}}\right)\leq\frac{C}{n_{1}}$ (D.10)
with $C=\max\{\frac{2c_{1}}{{\lambda_{\textup{min}}^{\infty}}^{2}}\sqrt{8n^{2}\|\Phi\|_{\infty}^{4}+2\|y\|^{4}},c_{2}\}$. This proves the claim for $\mathcal{X}$.
@@
By the formula for the derivative of the product and Cauchy-Schwarz’s inequality, the following estimate holds almost everywhere:
@@
Put $\overline{\lambda}=\min_{1\leq i\leq n}k_{\infty}(x,x_{i})$. The last summand can be further estimated with the first result in this theorem. There exists a positive constant $C$ not depending on $n_{1}$ such that:
$\frac{\partial}{\partial t}(y_{t}-G_{t}^{x})\leq\|k_{x\mathcal{X}}-k_{\infty}^{x}\|\|y-f_{t}\|+\sqrt{\frac{C}{n_{1}}}.$ (D.12)
Integrating we obtain:
$(y_{t}-G_{t}^{x})$ $\leq(y_{0}-G_{0}^{x})+\|k_{x\mathcal{X}}-k_{\infty}^{x}\|\|y-f_{0}\|t+\sqrt{\frac{C}{n_{1}}}t.$ (D.13)
we can finish by applying the elementary inequality $(a+b+c)^{2}\leq 3a^{2}+3b^{2}+3c^{2}$ for $a,b,c\geq 1$ and Hölder’s inequality. After that, Theorem B.3, Proposition B.4 and Lemma B.8 can be applied in the same fashion as in the proof of the training case, yielding positive constants $d_{1}$ and $d_{2}$ such that:
@@
$\leq 3\mathbb{E}[|y_{0}-G_{0}^{x}|^{2}]+3\mathbb{E}[\|k_{x\mathcal{X}}-k_{\infty}^{x}\|^{4}]^{\frac{1}{2}}\mathbb{E}[\|y-f_{0}\|^{4}]^{\frac{1}{2}}t^{2}+\frac{3Ct^{2}}{n_{1}}$ (D.15)
$\leq\frac{3d_{1}}{n_{1}}+\frac{6d_{2}}{n_{1}}\sqrt{8n^{2}\|\Phi\|_{\infty}^{4}+2\|y\|^{4}}t^{2}+\frac{3Ct^{2}}{n_{1}}$ (D.16)
$\leq C^{\prime}(1+t^{2})\frac{1}{n_{1}},$ (D.17)
with $C^{\prime}=\max\{3d_{1},6d_{2}\sqrt{8n^{2}\|\Phi\|_{\infty}^{4}+2\|y\|^{4}},3C\}$. Note that $C$ and $C^{\prime}$ do not depend neither on $t$ nor $n_{1}$. This finishes the proof.